Hello,
I am needing help on this money/percentage problem....I dont know where to start or how to set the equation up...If you could help me find the solution to this problem your help would be greatly appreciated.
A retired couple earned $656 last year from investments made in tax-free bonds. They invested some of the money at 8% and the rest at 10%. The total amount they invested was $7200. If the couple invested the $7200 for exactly one year, how much did they invest at teach interest rate? I have to identify each amount invested at each rate....
Thank you,
Mrs. Gibson

Answer by Earlsdon(6103)

Ok Mrs. Gibson,
First, let $x = the amount invested at 8% and the rest ($7200-$x) invested at 10%
The amount of interest earned on each of these amounts can be expressed, after changing the percentages to their decimal equivalents, as:
$x(0.08) and ($7200-$x)(0.1) and, the problem states the the sum of these two amounts of interest is $656.
So now you have enough to set up an equation in x.
0.08x + 0.1(7200-x) = 656 Simplify and solve for x.
0.08x + 720 - 0.1x = 656 Combine like-terms.
-0.02x + 720 = 656 Subtract 720 from both sides of the equation.
-0.02x = -64 Divide both sides by -0.02
x = 3200
So, $3200 was invested at 8% while $4000 ($7200-$3200) was invested at 10%